OpenStudy (anonymous):

1/cosx - cosx = sin x tan x prove the identity

OpenStudy (lgbasallote):

HI LANA!!! anyway...change tanx to \(\large \frac{\sin x}{\cos x}\) that should help

OpenStudy (lgbasallote):

and know that \[\sin^2 x = 1 - \cos^2 x\]

OpenStudy (ash2326):

\[\frac{1}{\cos x}-\cos x\] Take cos x common from both the terms \[\cos x(\frac{1}{\cos^2 x}-1)\] \[\cos x(\sec ^2 x-1)\] We know \[\sec^2 x-1=\tan^2 x\] so \[\cos x\times \tan^2 x\] Can you take it from here?

OpenStudy (lalaly):

\[\frac{1}{cosx}-\frac{\cos^2x}{cosx}=\frac{\sin^2x}{cosx}=\frac{sinx}{cosx}\times sinx\]

OpenStudy (lgbasallote):

hello @lalaly :p btw @alexeis_nicole the solution i said was for right hand side

OpenStudy (anonymous):

@ lgbasallote yep. I'm doing that right now. but i got 1/cosx ... i'm trying to see where i went wrong for it

OpenStudy (lgbasallote):

\[\sin x \tan x = \sin x (\frac{\sin x}{\cos x}) = \frac{\sin^2 x}{\cos x}\] \[\frac{1 - \cos^2 x}{\cos x} = \frac{1}{\cos x} - \frac{\cos^2 x}{\cos x}\]

OpenStudy (anonymous):

oh ! @lgbasallote okay, sorry i was kinda trying to figure out how you did that. now i get it. thank you !